OpenStudy (twizttiez):

Which is an equation in point-slope form for the given point and slope? Point: (2,-6) ; Slope:-3/4 A.y-6=-3/4(x+2) B.y-6=3/4(x-2) C.y+6=-3/4(x-2) D.y+6=-3/4(x+2)

OpenStudy (twizttiez):

I think it's C but not 100% sure

OpenStudy (anonymous):

yes, it is c.

OpenStudy (anonymous):

Yes it's C

OpenStudy (twizttiez):

Ok can i get some help on some more?

OpenStudy (anonymous):

yes.

OpenStudy (twizttiez):

Thanks!

OpenStudy (anonymous):

\[y-y _{1} = m(x-x _{1})\]

OpenStudy (dmndlife24):

Using the formula y - y1 = m (x - x1) 2 = x1 -6 = y1 y + 6 = -3/4(x - 2) The answer is C. :D

OpenStudy (dmndlife24):

m = slope = -3/4

OpenStudy (twizttiez):

What is the point-slope form of the equation of a line that passes through the point (3, 5) and has a slope of 2? A.y+3=2(x+5) B.y-3=2(x-5) C.y-5=2(x-3) D.y-5=-2(x-3)

OpenStudy (twizttiez):

@dmndlife24 @silverr @dimensionx

OpenStudy (dmndlife24):

Put it in a new post...one question per post

OpenStudy (twizttiez):

Ok